OpenStudy (anonymous):

How do I solve this word problem? The perimeter of a rectangle is 160 ft. One fourth the length is the same as twice the width. Find the dimensions of the rectangle.

OpenStudy (anonymous):

consider the length as X and the width as Y . Substitute them for both conditions to get simultaneous linear equations :)

OpenStudy (anonymous):

.25L = 2W Start by dividing both sides by .25

OpenStudy (anonymous):

L = 8W?

OpenStudy (anonymous):

yes you're right there

OpenStudy (anonymous):

now apply that in the formula for the perimeter

OpenStudy (anonymous):

Correct! So L=1 unit and W=8 units. So if A =L*W, then L=A/W

OpenStudy (anonymous):

You don't have the area given. You have the perimeter .. the perimeter P=2L+2W

OpenStudy (anonymous):

You're right. Hold on, I'm being distracted to your detriment, and I apologize.

OpenStudy (anonymous):

wait im confused.

OpenStudy (anonymous):

yes L=8W is correct

OpenStudy (anonymous):

now perimeter P = 160 = 2L + 2W .. use L=8W and and find L . Then find W

OpenStudy (anonymous):

Okay, I'm back. Now we know that L is 8 units and W is 1 unit, so the sum of 2 sides must be 9 units. It follows that all 4 sides must be 18 units. So dividing the perimeter by 18 gives you the length of 1 unit, and caution: It is NOT a nice neat number! :-) So the W will be equal to 1 unit and the L will be equal to 8 units.